Jensen Associates provides marketing services for a number of small manufacturing firms. Jensen receives a commission of 10 percent of sales. Operating costs are as follows:

Unit-level costs $ 0.04 per sales dollar
Sales-level costs $ 300 per sales order
Customer-level costs $ 900 per customer per year
Facility-level costs $ 60,000 per year

(a) Determine the minimum order size in sales dollars for Jensen to break even on an order.

(b) Assuming an average customer places four orders per year, determine the minimum annual sales required to break even on a customer.

(c) What is the average order size in (b)?

(d) Assuming Jensen currently serves 100 customers, with each placing an average of four orders per year, determine the minimum annual sales required to break even.

(e) What is the average order size in (d)?
Answer #1

Solution :

a. minimum order size in sales dollars for Jensen to break even on an order :

                Sales level cost / contribution margin ratio = 300 / 0.06* = 5000.

                *(0.10 - 0.04)

b. minimum annual sales required to break even on a customer. :

[(sales level cost X Avg customer order per year) + customer level cost] / Contribution margin ratio

= [( 300 x 4) + 900 ] / 0.06

= 35000

c. average order size in (b) = 35000 / 4 = 8750

d. Assuming Jensen currently serves 100 customers, with each placing an average of four orders per year, the minimum annual sales required to break even wil be as follows :

PARTICULARS

order level cost ( 300 X 4 X 100)

120000

customer level cost (900 X 100)

90000

facility level cost

60000

Total cost

270000

Divide by Contribution margin ratio

0.06

minimum annual sales required to break even

4500000

e. average order size in (d) = 4500000 / (100 x 4 ) = 11250